OpenStudy (anonymous):

The sum of three consecutive odd integers is 57. What are the three numbers?

OpenStudy (anonymous):

let the 1st integer be x then 2nd odd integer will be x+2 and third odd integer will be x+4 sp x+x+2+x+4 =57 3x=51 x=17 1st odd integer =17 2nd odd integer will be x+2 =17+2 =19 and third odd integer will be x+4 = 17+4 =21

OpenStudy (anonymous):

so 17,19,21 cross check 17+19+21= 57